[Brahms-l] Shift report 20050427 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Apr 27 2005 - 00:14:48 EDT
Shifter leader was FR
Shifters were KH


 - 1 store from 6:15 pm; still in progress. Will be kept until 2:00 am.
 - Data with FS @ 8deg, 2442A.
   5 runs 14294-14298, each ~1 hour, with a total of 5350 trig2 events.
   Run 14299 : in progress
 - No detector problem.


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Apr 27 00:15:02 2005

This archive was generated by hypermail 2.1.8 : Wed Apr 27 2005 - 00:15:18 EDT